Given that the cost of sales of a company XYZ for the year 2005 was
ksh 570,000 and the average stock amount to ksh 124,250.
Calculate the stock turnover
1
Expert's answer
2018-10-30T16:35:09-0400

The stock turnover is calculated as

Stock turnover=(Cost of sales)/(Average stock held)=(570,000)/(124,250)=4.59 times